# Emergent symmetry in a two-Higgs-doublet model from quantum information and magic

### v1

@@ Abstract
Studies of scattering processes in scalar models with two Higgs doublets have recently hinted at a connection between the absence of flavour-space entanglement in $\Phi^{+}\Phi^{0}$ scattering and an emergent $\mathrm{SO}(8)$ symmetry in the scalar potential. We extend the analysis to all scattering channels with two particles in the external states by treating the process as a four-qubit system in the weak isospin and flavour subspaces of the $2$-particle state. We work with a generic quantum information-theoretic principle encoded by the commutativity of the initial state density matrix with the transition matrix (at leading order in perturbation theory). This yields a special case of the entanglement minimisation conditions previously derived in the literature, and we interpret the principle in terms of the conservation of non-stabiliserness (or “magic”). Working at leading order in the quartic couplings, we find a consistent set of conditions that implies an $\mathrm{SO}(8)$ symmetry on the quartic part of the potential for scattering an arbitrary initial state, but a smaller $\mathrm{SU}(2)_{R}$ symmetry when the initial state is chosen to have definite isospin. This follows by accounting for Bose symmetry in the initial state, which introduces entanglement between the isospin and flavour subspaces.
@@ I Introduction
Our aim here is to elucidate the emergence of symmetry in the 2HDM from scalar $2\to 2$ scattering processes, focusing on initial states with two incoming particles (i.e., total particle number $\mathcal{P}=2$), while accounting for Bose symmetry and $\mathrm{SU}(2)_{L}$ invariance. We show that the emergence of the maximal $\mathrm{SO}(8)$ symmetry in the case of $\Phi^{+}\Phi^{0}$ scattering can be framed as following from the condition that the initial and final state density matrices are equal up to first order in the quartic couplings of the theory. This is a stronger condition than requiring that the transition matrix $\mathcal{T}$ have vanishing entanglement power Zanardi et al. (2000), and we show that it is equivalent to requiring the conservation of non-stabiliserness, colloquially referred to as magic White and White (2024), which characterises those states of a system that confer an advantage in quantum computing. Concretely, our condition is equivalent to the requirement that the initial-state density matrix commutes with $\mathcal{T}$.
@@
This constraint also restricts the form of the $\mathcal{T}$-matrix to be proportional to the identity gate for an arbitrary initial density matrix. This is similar to the case of minimally entangling $np$ scattering, [^1] An important difference is that $\mathcal{T}$ can also access the $\mathrm{SWAP}$ gate in $np$ scattering. where the condition follows from rotational invariance in $s$-wave scattering Low and Mehen (2021); Liu et al. (2023). We note however that Bose symmetry plays an interesting role in restricting the form of $\rho$ and therefore allowing more flexibility in the structure of the $\mathcal{T}$-matrix. We will discuss below that the maximal $\mathrm{SO}(8)$ symmetry in the 2HDM potential follows from the fact that $|\Phi^{+}\Phi^{0}\rangle$ is a superposition of the triplet and singlet states of $\mathrm{SU}(2)_{L}$ with vanishing projection of the third component of weak isospin. Our main aim is to extend the study of the emergence of symmetry to all $2\to 2$ scattering processes with non-vanishing particle number within the same explicitly isospin-invariant formalism. We find that the scattering of an arbitrary initial state results in an $\mathrm{SO}(8)$ symmetry acting on the quartic part of the potential. Interestingly, when Bose and $\mathrm{SU}(2)_{L}$ symmetry are accounted for in the initial state, consistent conditions imprinting an $\mathrm{SU}(2)_{R}$ symmetry onto the scalar potential can be derived after imposing $[\mathcal{T},\rho]=0$ when scattering states of definite weak isospin. As we will see, this analysis complements the existing entanglement results regarding the emergence of the alignment limit (which ensures a SM-like Higgs boson). Recent related work can be found in the study of magic in the neutrino sector in Ref. Chernyshev et al. (2024), in the context of QED in Ref. Liu et al. (2025), and investigations of the magic generated in top quark pair production at the Large Hadron Collider in Refs. White and White (2024); CMS Collaboration (2025); Aoude et al. (2025).
@@
This paper is structured as follows. In Sec. II, we provide a convenient introduction to the concept of non-stabiliserness, a brief review of the salient features of the 2HDM and an explanation of how to treat scalar scattering in flavour-space as a two qubit system. In Sec. III, we compute the initial and final state magic for the process $\Phi^{+}\Phi^{0}\to\Phi^{+}\Phi^{0}$ at leading order in perturbation theory, and demonstrate the symmetries that emerge from conservation of magic. In Sec IV, we extend our formalism to cover all $2\to 2$ scalar scattering processes involving states with total particle number $\mathcal{P}=2$ and investigate the resulting emergent symmetries. We conclude in Sec. V.
@@ II.1 Magic in Quantum Information Theory
where $i_{k}=0,1,2,3$. These Pauli strings generate the Pauli group, essentially consisting of the Pauli strings weighted by phases of ±$\pm 1,\pm i$. Then one may define the $n$-qubit Clifford group $\mathbb{C}_{n}$ consisting of all unitary operators fixing the Pauli group; that is,
@@
The physical significance of these concepts is this: The Gottesman-Knill theorem Gottesman (1998), by now a standard result in quantum information theory, asserts that quantum circuits comprised of Clifford gates may be efficiently simulated on a classical computer, even though they may produce highly entangled states. In other words, entanglement alone is insufficient to guarantee computational advantage; an additional resource is required. This resource is referred to as non-stabiliserness or magic. Magic states are precisely those states that are not stabiliser states.
@@
[…]
@@ II.3 Scalar Scattering of Qubits
×$\quad\times\delta^{(3)}(\mathbf{p}_{1}-\mathbf{q}_{1})\delta^{(3)}(\mathbf{p}_{2}-\mathbf{q}_{2})\ ,$ (9)
@@ III Magic Conservation and the Alignment Limit
and hence $\rho_{f}=\rho_{0}+\mathcal{O}(\lambda^{2})$. In fact, it is straightforward to verify that Eq. (15) is the necessary condition for $\langle ij|[\mathcal{T},\rho_{0}]|kl\rangle=0+\mathcal{O}(\lambda^{2})$; thus, conservation of magic is equivalent to the vanishing of the commutator $[\mathcal{T},\rho_{0}]$. This is in contrast to other information-theoretic measures, such as the fidelity or von Neumann entropy, which are trivially conserved at leading order, and thus do not lead to an interesting constraint. On the other hand, the vanishing of this commutator will surely entail the leading-order conservation of various other information-theoretic functions; thus, if one were attempting to formulate a generalised quantum-information principle applicable to a host of measures, one might be naturally inclined to this scenario. We hereafter treat the commutativity of $\rho_{0}$ and $\mathcal{T}$ as our guiding constraint.
@@ IV.1 Generalised Scattering Formalism
$\mathbf{C}^{0,0}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}0&1\\ -1&0\end{pmatrix}\ ,$ (25)
$\mathbf{C}^{1,m}$ =$=\begin{cases}\begin{pmatrix}1&0\\ 0&0\end{pmatrix}&m=1\\ \frac{1}{\sqrt{2}}\begin{pmatrix}0&1\\ 1&0\end{pmatrix}&m=0\\ \begin{pmatrix}0&0\\ 0&1\end{pmatrix}&m=-1\end{cases}\ .$ (26)
@@
$\bm{\lambda}_{T}$ $=\begin{pmatrix}\lambda_{1}&\tfrac{1}{\sqrt{2}}\lambda_{6}&\tfrac{1}{2}\lambda_{5}\\ \tfrac{1}{\sqrt{2}}\lambda_{6}^{*}&\tfrac{1}{2}(\lambda_{3}+\lambda_{4})&\tfrac{1}{\sqrt{2}}\lambda_{7}\\ \tfrac{1}{2}\lambda_{5}^{*}&\tfrac{1}{\sqrt{2}}\lambda_{7}^{*}&\lambda_{2}\end{pmatrix}\ ,$
@@ IV.2 Maximal symmetry from $\Phi\Phi\to\Phi\Phi$ scattering
[…]
$\begin{pmatrix}|\alpha_{0,0}|^{2}|\beta_{0,0}|^{2}&\alpha_{0,0}\alpha_{1,m^{\prime}}^{*}\beta_{0,0}\beta_{1,n^{\prime}}^{*}\\ \alpha_{1,m}\alpha_{0,0}^{*}\beta_{1,n}\beta_{0,0}^{*}&\alpha_{1,m}\alpha_{1,m^{\prime}}^{*}\beta_{1,n}\beta_{1,n^{\prime}}^{*}\end{pmatrix}\Bigg{]}\ .$
@@ Acknowledgements
We thank Rafael Aoude, Hannah Banks, Tomas L. Howson, Raymond R. Volkas and Chris D. White for useful conversations and comments on the draft. All authors are supported by the Australian Research Council grant CE200100008. MJW is further supported by the Australian Research Council grant DP220100007. EW acknowledges the support he has received for this research through the provision of an Australian Government Research Training Program Scholarship.

### v2

# Emergent symmetry in a two-Higgs-doublet model from quantum information and nonstabiliserness

@@ Abstract
Studies of scattering processes in scalar models with two Higgs doublets have recently hinted at a connection between the absence of flavour-space entanglement in $\Phi^{+}\Phi^{0}$ scattering and an emergent $\mathrm{SO}(8)$ symmetry in the scalar potential. We extend the analysis to all scattering channels with two particles in the external states by treating the process as a four-qubit system in the weak isospin and flavour subspaces of the $2$-particle state. We work with a generic quantum information-theoretic principle encoded by the commutativity of the initial state density matrix with the transition matrix (at leading order in perturbation theory). This yields a special case of the entanglement minimisation conditions previously derived in the literature, and we interpret the principle in terms of the conservation of nonstabiliserness (or “magic”). Working at leading order in the quartic couplings, we find a consistent set of conditions that implies an $\mathrm{SO}(8)$ symmetry on the quartic part of the potential for scattering an arbitrary initial state, but a smaller $\mathrm{SU}(2)_{R}$ symmetry when the initial state is chosen to have definite isospin. This follows by accounting for Bose symmetry in the initial state, which introduces entanglement between the isospin and flavour subspaces.
@@ I Introduction
Our aim here is to elucidate the emergence of symmetry in the 2HDM from scalar $2\to 2$ scattering processes, focusing on initial states with two incoming particles (i.e., total particle number $\mathcal{P}=2$), while accounting for Bose symmetry and $\mathrm{SU}(2)_{L}$ invariance. We show that the emergence of the maximal $\mathrm{SO}(8)$ symmetry in the case of $\Phi^{+}\Phi^{0}$ scattering can be framed as following from the condition that the initial and final state density matrices are equal up to first order in the quartic couplings of the theory. This is a stronger condition than requiring that the transition matrix $\mathcal{T}$ have vanishing entanglement power Zanardi et al. (2000), and we show that it is equivalent to requiring the conservation of nonstabiliserness, colloquially referred to as magic White and White (2024), which characterises those states of a system that confer an advantage in quantum computing. Concretely, our condition is equivalent to the requirement that the initial-state density matrix commutes with $\mathcal{T}$.
@@
This constraint also restricts the form of the $\mathcal{T}$-matrix to be proportional to the identity gate for an arbitrary initial density matrix. This is similar to the case of minimally entangling $np$ scattering, [^1] An important difference is that $\mathcal{T}$ can also access the $\mathrm{SWAP}$ gate in $np$ scattering. where the condition follows from rotational invariance in $s$-wave scattering Low and Mehen (2021); Liu et al. (2023). We note however that Bose symmetry plays an interesting role in restricting the form of $\rho$ and therefore allowing more flexibility in the structure of the $\mathcal{T}$-matrix. We will discuss below that the maximal $\mathrm{SO}(8)$ symmetry in the 2HDM potential follows from the fact that $|\Phi^{+}\Phi^{0}\rangle$ is a superposition of the triplet and singlet states of $\mathrm{SU}(2)_{L}$ with vanishing projection of the third component of weak isospin. Our main aim is to extend the study of the emergence of symmetry to all $2\to 2$ scattering processes with non-vanishing particle number within the same explicitly isospin-invariant formalism. We find that the scattering of an arbitrary initial state results in an $\mathrm{SO}(8)$ symmetry acting on the quartic part of the potential. Interestingly, when Bose and $\mathrm{SU}(2)_{L}$ symmetry are accounted for in the initial state, consistent conditions imprinting an $\mathrm{SU}(2)_{R}$ symmetry onto the scalar potential can be derived after imposing $\mathcal{T},\rho]=0$ when scattering states of definite weak isospin. As we will see, this analysis complements the existing entanglement results regarding the emergence of the alignment limit (which ensures a SM-like Higgs boson). Recent related work can be found in the study of magic in the neutrino sector in Ref. Chernyshev et al. (2024), in the context of QED in Ref. Liu et al. (2025), and investigations of the magic generated in top quark pair production at the Large Hadron Collider in Refs. White and White (2024); CMS Collaboration (2025); Aoude et al. (2025).
@@
This paper is structured as follows. In Sec. II, we provide a convenient introduction to the concept of nonstabiliserness, a brief review of the salient features of the 2HDM and an explanation of how to treat scalar scattering in flavour-space as a two qubit system. In Sec. III, we compute the initial and final state magic for the process $\Phi^{+}\Phi^{0}\to\Phi^{+}\Phi^{0}$ at leading order in perturbation theory, and demonstrate the symmetries that emerge from conservation of magic. In Sec IV, we extend our formalism to cover all $2\to 2$ scalar scattering processes involving states with total particle number $\mathcal{P}=2$ and investigate the resulting emergent symmetries. We conclude in Sec. V.
@@ II.1 Magic in Quantum Information Theory
where $i_{k}=0,1,2,3$. These Pauli strings generate the Pauli group, essentially consisting of the Pauli strings weighted by phases of $\pm 1,\pm i$. Then one may define the $n$-qubit Clifford group $\mathbb{C}_{n}$ consisting of all unitary operators fixing the Pauli group; that is,
@@
The physical significance of these concepts is this: The Gottesman-Knill theorem Gottesman (1998), by now a standard result in quantum information theory, asserts that quantum circuits comprised of Clifford gates may be efficiently simulated on a classical computer, even though they may produce highly entangled states. In other words, entanglement alone is insufficient to guarantee computational advantage; an additional resource is required. This resource is referred to as nonstabiliserness or magic. Magic states are precisely those states that are not stabiliser states.
@@
$+\lambda_{4}(\Phi_{1}^{\dagger}\Phi_{2})(\Phi_{2}^{\dagger}\Phi_{1})+\left(\frac{\lambda_{5}}{2}(\Phi_{1}^{\dagger}\Phi_{2})^{2}+\lambda_{6}(\Phi_{1}^{\dagger}\Phi_{1})(\Phi_{1}^{\dagger}\Phi_{2})+\lambda_{7}(\Phi_{2}^{\dagger}\Phi_{2})(\Phi_{1}^{\dagger}\Phi_{2})+\text{h.c.}\right).$ (5)
@@ II.3 Scalar Scattering of Qubits
$\quad\times\delta^{(3)}(\mathbf{p}_{1}-\mathbf{q}_{1})\delta^{(3)}(\mathbf{p}_{2}-\mathbf{q}_{2})\ ,$ (9)
@@ III Magic Conservation and the Alignment Limit
and hence $\rho_{f}=\rho_{0}+\mathcal{O}(\lambda^{2})$. In fact, it is straightforward to verify that Eq. (15) is the necessary condition for $\langle ij|[\mathcal{T},\rho_{0}]|kl\rangle=0+\mathcal{O}(\lambda^{2})$; thus, conservation of magic is equivalent to the vanishing of the commutator $\mathcal{T},\rho_{0}]$. This is in contrast to other information-theoretic measures, such as the fidelity or von Neumann entropy, which are trivially conserved at leading order, and thus do not lead to an interesting constraint. On the other hand, the vanishing of this commutator will surely entail the leading-order conservation of various other information-theoretic functions; thus, if one were attempting to formulate a generalised quantum-information principle applicable to a host of measures, one might be naturally inclined to this scenario. We hereafter treat the commutativity of $\rho_{0}$ and $\mathcal{T}$ as our guiding constraint.
@@ IV.1 Generalised Scattering Formalism
$\mathbf{C}^{0,0}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}0&1\\ -1&0\end{pmatrix}\$, (25)
$\mathbf{C}^{1,m}$ $=\begin{cases}\begin{pmatrix}1&0\\ 0&0\end{pmatrix}&m=1\\ \frac{1}{\sqrt{2}}\begin{pmatrix}0&1\\ 1&0\end{pmatrix}&m=0\\ \begin{pmatrix}0&0\\ 0&1\end{pmatrix}&m=-1\end{cases}\ .$ (26)
@@
$\bm{\lambda}_{T}$ $=\begin{pmatrix}\lambda_{1}&\tfrac{1}{\sqrt{2}}\lambda_{6}&\tfrac{1}{2}\lambda_{5}\\ \tfrac{1}{\sqrt{2}}\lambda_{6}^{*}&\tfrac{1}{2}(\lambda_{3}+\lambda_{4})&\tfrac{1}{\sqrt{2}}\lambda_{7}\\ \tfrac{1}{2}\lambda_{5}^{*}&\tfrac{1}{\sqrt{2}}\lambda_{7}^{*}&\lambda_{2}\end{pmatrix}\$,
@@ IV.2 Maximal symmetry from $\Phi\Phi\to\Phi\Phi$ scattering
$\Bigg{[}$ $\begin{pmatrix}\lambda_{S}&0\\ 0&(\lambda_{T})_{n,n^{\prime}}\delta_{m,m^{\prime}}\\ \end{pmatrix}$, (34)
$\begin{pmatrix}|\alpha_{0,0}|^{2}|\beta_{0,0}|^{2}&\alpha_{0,0}\alpha_{1,m^{\prime}}^{*}\beta_{0,0}\beta_{1,n^{\prime}}^{*}\\ \alpha_{1,m}\alpha_{0,0}^{*}\beta_{1,n}\beta_{0,0}^{*}&\alpha_{1,m}\alpha_{1,m^{\prime}}^{*}\beta_{1,n}\beta_{1,n^{\prime}}^{*}\end{pmatrix}\Bigg{]}\$.
@@ V Conclusions
We comment here on the consistency of our results with previous studies. Even though magic conservation is more constraining than entanglement minimisation, the results of Ref. Chang and Jacobo (2024) are not inconsistent with our results because we do not consider the same class of scattering processes; namely, we concentrate only on $\mathcal{P}=2$ channels. We anticipate that the inclusion of isospin as an additional quantum number may lead to different results even in the $\mathcal{P}=0$ channels too. Ref. Kowalska and Sessolo (2024) studies other $\mathcal{P}=2$ channels in the context of entanglement minimisation, but their results aggregate over all scattering processes; further, we find that symmetrisation of the initial state (for the case of identical particles) plays an important role in our analysis.
@@ Acknowledgements
We thank Rafael Aoude, Hannah Banks, Tomas L. Howson, Raymond R. Volkas and Chris D. White for useful conversations and comments on the draft. All authors are supported by the Australian Research Council grant CE200100008. MJW is further supported by the Australian Research Council grant DP220100007. ENVW acknowledges the support he has received for this research through the provision of an Australian Government Research Training Program Scholarship.